# Multimodal atmospheric super-resolution with deep generative models

### v1

@@ 4.3 Multimodal super-resolution
In this section, we evaluate one of the greatest strengths of zero-shot sampling from generative models, namely, their ability to assimilate multiple sources of data during test time. Specifically, we perform experiments that rely on observations of both the low-resolution ERA5 observations on a coarse and structured grid, as well as IGRA observations on the unstructured data (previously introduced as the sample-IGRA+LR configuration). In Figure 10 we show the spectral recovery properties of the multifidelity assimilation versus those obtained solely from IGRA and the structured grid observations. One of the key benefits of using data from the low-resolution ERA5 in addition to IGRA is the significant reduction in uncertainty in the lower wavenumbers for the wind-speed spectra in addition to a correction of the high-wavenumber errors obtained when solely using IGRA. This is also observed in longitudinal traces in Figure 12 where the variance in samples is reduced given more relevant information. When assessing the scatter of the generated values for exemplar state recovery at the IGRA sensor locations, as shown in Figure 11, we observe that the generative model is able to sample in a manner that is closer to the observed IGRA magnitudes despite using both low-resolution ERA5 and IGRA for score-matching. This indicates an ability to balance the influence of various sources of data based on the proximity to their observation locations. This is reinforced by evidence in Figure 7, where scatter plots for reconstructions at non-IGRA sensing locations indicate an improved reconstruction of ERA5 now that coarse-grid ERA5 observations are provided during zero-shot score matching. We remark that while using low-resolution observations of reanalysis represents an unrealistic source of data during real-time data fusion, it represents a proof-of-concept of fusing another high-quality source of data to reduce uncertainties during assimilation.
@@ 5 Discussion and Conclusion
Rapid assimilation of observations from real-time platforms has the potential to provide significant gains in the accurate assessment of high-dimensional dynamical systems. The motivation of this study was to propose avenues for reducing the significant computational costs for achieving the same using classical numerical methods for data assimilation of high-dimensional multiscale systems. To that end, this study proposed the construction of an unconditional generative model that is trained to match the score of the underlying probability distribution from which arbitrary training data may be sampled. Once this score is approximated with a neural network, a zero-shot score-matching technique built on a Bayesian formulation can be used to rapidly assimilate sparse multifidelity observational data with negligible real-time costs. In this study, we construct an experiment to recover the full state of a reanalysis dataset given partial observations from a coarse-grid observation of the same as well as when sparse and spatially unstructured data are obtained from a different dataset altogether. In both such examples, we observe that the score-matching technique is proficient in recovering the original state of the reanalysis data on the fine grid. Notably, when both sets of partial observations are used for a multifidelity state recovery exercise, we observe that the data fusion strategy is able to balance proximity to both data sets based on their spatial sensing locations. This indicates that zero-shot score matching can be used as an effective and computationally inexpensive tool to accurately perform data fusion from data of varying sources and spatiotemporal fidelity. Based on the promising results from this study, future extensions of this framework include the integration of dynamical considerations, i.e., performing state vector trajectory reconstructions given time-series traces at observation locations, integrating both Eulerian and Lagragian observations within a common fusion paradigm, and performing optimal sensor placement for high quality state recovery.
@@ Acknowledgments
This research used resources of the Argonne Leadership Computing Facility, which is a U.S. Department of Energy Office of Science User Facility operated under contract DE-AC02-06CH11357. DC, HG, and RM acknowledge support from DOE ASCR award “Inertial neural surrogates for stable dynamical prediction" and DOE FES award “DeepFusion Accelerator for Fusion Energy Sciences in Disruption Mitigation". DC, HG, RM also acknowledge the support of computing resources through an AI For Science allocation at the National Energy Research Scientific Computing Center (NERSC) and from Penn State Institute for Computational and Data Sciences.

### v2

@@ 4.3 Multimodal super-resolution
In this section, we evaluate one of the greatest strengths of zero-shot sampling from generative models, namely, their ability to assimilate multiple sources of atmospheric data during test time. Therefore, we denote these studies as multimodal super-resolution. Note that modality refers to platforms of measurement as against different data types. First, we perform experiments that rely on observations of both the low-resolution ERA5 observations on a coarse and structured grid, as well as IGRA observations on the unstructured data (previously introduced as the sample-IGRA+LR configuration). In Figure 12 we show the spectral recovery properties of the multifidelity assimilation versus those obtained solely from IGRA and the structured grid observations. One of the key benefits of using data from the low-resolution ERA5 in addition to IGRA is the significant reduction in uncertainty in the lower wavenumbers for the wind-speed spectra in addition to a correction of the high-wavenumber errors obtained when solely using IGRA. This is also observed in longitudinal traces in Figure 14 where the variance in samples is reduced given more relevant information. When assessing the scatter of the generated values for exemplar state recovery at the IGRA sensor locations, as shown in Figure 13, we observe that the generative model is able to sample in a manner that is closer to the observed IGRA magnitudes despite using both low-resolution ERA5 and IGRA for score-matching. This indicates an ability to balance the influence of various sources of data based on the proximity to their observation locations. This is reinforced by evidence in Figure 15, where scatter plots for reconstructions at non-IGRA sensing locations indicate an improved reconstruction of ERA5 now that coarse-grid ERA5 observations are provided during zero-shot score matching. We remark that while using low-resolution observations of reanalysis represents an unrealistic source of data during real-time data fusion, it represents a proof-of-concept of fusing another high-quality source of data to reduce uncertainties during assimilation.
@@
Next, we showcase the model’s ability to assimilate data coming from a dynamical core along with observations coming from a sparse and unstructured data set. For evaluation, LUCIE generates ensembles of 7-day forecasts, initialized at 12:00 UTC on randomly sampled days in 2020, serving as measurements for the EDM model to perform posterior sampling. We also remind the reader that the LUCIE forecasts are provided on a much coarser grid than that of the training data. Data fusion that solely uses LUCIE observations during the sampling process are denoted ‘LUCIE’. Next, we also use measurements from IGRA dataset in a composite likelihood, as mentioned in Equation 14, to perform multimodal data fusion, denoted ‘IGRA+LUCIE’. We remark that these fusion implementations do not reinitialize the LUCIE forecast model. In other words, we simply assimilate precomputed forecasts from a LUCIE deployment on the same initial condition. The hyperparameter $\lambda$ in Equation 14, adapts the relative influence of these two datasets. Table 2 shows the existence of an ideal range of $\lambda$ to obtain the optimal RMSE.
(a) IGRA+LUCIE
(b) IGRA+LUCIE Reinitialized
Figure 17: Comparison of IGRA+LUCIE (a) and IGRA+LUCIE Reinitialized (b). In (a), sequences of LUCIE forecasts and measurements from IGRA are fused for super-resolving at each time step independently. In (b), a LUCIE forecast is fused with IGRA measurements and super-resolved, downsampled, and reinitialized for forecast iteratively.
Figure 18: RMSE over 7 days for selected variables. The red dashed line denotes the unconditional baseline that is averaged over 28 random samples. The rest are computed as the RMSE of averaged 16 ensemble members with different initial conditions.
$\lambda$ Day 1 Day 3 Day 5 Day 7
0.0 1.934 3.160 4.273 4.938
0.5 1.678 2.828 3.919 4.241
1.0 1.674 2.793 3.849 4.166
2.0 1.599 2.584 3.574 3.825
4.0 1.661 2.677 3.625 3.910
Table 2: 2m temperature RMSE by lead time for different $\lambda$ (Equation 14).
Figure  18 shows that sampling through observations of precomputed LUCIE forecasts initially achieves low error but degrades with lead time due to the onset of deterministic chaos. This is an expected trend - we emphasize that observations are not being used to correct our dynamical core, as is common in conventional data assimilation, but a previous forecast from an uncorrected dynamical core is being used to guide the sampling of the posterior of our diffusion model. A true equivalence with data assimilation would need the utilization of these reconstructions as fresh initial conditions of the dynamical core (which we will return to shortly). Next, as observed in Section 4.2, sampling with IGRA gives higher accuracy in locations where observations are abundantly available. This causes it to have a relatively constant RMSE but one that is significantly lower than unconditional sampling from the prior. The combined likelihood from the IGRA+LUCIE approach yields improved performance in the sampling, starting closer to LUCIE and showing significant improvement at larger lead time. However, we observe, as shown for when only LUCIE was used in the data fusion process, that growth in emulator errors ultimately causes the reconstructions to have higher errors than merely utilizing IGRA data at each snapshot. We also remark that the fusion step is performed in lock-step with IGRA observations, which are available every 12 hours despite LUCIE being available every 6 hours.
Finally, we introduce a data fusion formulation that implements a reinitialization for LUCIE for each timestep of a super-resolution being performed sequentially in time, denoted ‘IGRA+LUCIE Reinitialized’. Figure 17 represents a schematic that contrasts the previous approach to incorporating LUCIE forecasts, versus what is now proposed, where the super-resolution via diffusion sampling is used to reinitialize the LUCIE dynamical core for a one step forecast. This setup represents a scenario that is analogous to classical atmospheric data assimilation, with the important caveat that the covariance matrix is pre-specified for rapid inverse computation. In Figure 19, we observe that the proposed approach leads to improved or similar performance later in the forecast horizon when compared to solely using precomputed LUCIE or IGRA+LUCIE forecasts. In terms of sampling times mentioned in Table 3, we noticed the inclusion of IGRA increases the sampling time. However, we do not detect any major differences between sampling times of IGRA, IGRA+LUCIE and IGRA+LUCIE Reinitialized. We hypothesize further improved gains, at the cost of reduced computational efficiency, if covariances are updated and inverted for each reinitialization step, but we leave this for a future work due to its computational intractability.
Method Mean Time (s) Std (s)
LUCIE 1.995 0.00007
IGRA 2.298 0.00504
IGRA+LUCIE 2.298 0.00073
IGRA+LUCIE Reinitialized 2.293 0.00061
Table 3: Sampling run times per ensemble member (mean and standard deviations) for different posterior sampling methods.
The RMSE of an unconditional model and the low-resolution ERA5 serve as the the two extremes for our assessments: the unconditional being the worst case and the low-resolution ERA5 providing the best reconstructions. Furthermore, to gauge the temporal consistency of our results, we plot the ensemble mean of 2m temperature at particular latitudes with respect to lead time in Figure 19. We observe at latitude 0.7 and longitudes from -120 to -180 that IGRA+LUCIE samples perfectly match the temporal trend with ERA5 as against reconstructions that solely utilize IGRA. This shows that AI-based emulators can provide valuable dynamical information for temporal consistency in independently reconstructed samples from sparse observations.
Figure 19: The time series of ensemble mean 2m temperature at 0.7 latitude with initial condition at 12:00 UTC January 1st 2020. IGRA+LUCIE Reinitialized is able to recover oscillatory trends in the temperature that are typically missed by IGRA-based reconstructions alone.
@@ 5 Discussion and Conclusion
Rapid fusion of observations from different, real-time platforms has the potential to provide significant gains in the accurate reconstruction of high-dimensional dynamical systems. The motivation of this study was to propose avenues for reducing the significant computational costs for achieving the same when using classical numerical methods for data assimilation. To that end, this study proposed the construction of an unconditional generative model that is trained to match the score of the underlying probability distribution from which arbitrary training data may be sampled. Once this score is approximated with a neural network, a zero-shot score-matching technique built on a Bayesian formulation can be used to rapidly assimilate sparse multifidelity observational data with negligible real-time costs. In this study, we first construct an experiment to recover the full state of a reanalysis dataset given partial observations from a coarse-grid observation of the same. This represents an assessment of the method’s ability to recover reanalysis when shown partial information. Next, we perform experiments to assess how the score-matching can recover reanalysis when shown information from a real-world radiosonde dataset as well as snapshots from an atmospheric emulator. These experiments indicate that our proposed method can balance the influence of the various observations during the reanalysis reconstruction process. Finally, we perform an experiment that is conceptually closer to real-world atmospheric data assimilation, where the instantaneous reconstruction of the atmosphere is re-utilized as an initial condition for the emulator forecast. These results show that improved gains in accuracy may be obtained when a dynamical core is reinitialized for a one-step forecast in a sequential data fusion task. Specifically, we note that the best practices of atmospheric data assimilation, such as improved covariance matrices can lead to significant gains on top of the proof-of-concept proposed here. However, we note that offline precomputation of the forecasts from a dynamical core also provide competitive results at much reduced computational costs. More importantly, to the best of our knowledge, this work represents a combined data assimilation and super-resolution task that allows for the use of observations from various dynamical models and observations on varying grids. When compared to adjoint-based methods, our approach does not require the computation of gradients through a multistep dynamical core roll-out. When compared with ensemble methods, our approach allows for the assimilation of dynamical predictions on a different grid from the final output which is that of the training data resolution.
In all experiments, we observe that the score-matching technique generates reconstructions for the atmosphere that outperform an unconditional sampling from the pretrained diffusion model indicating an information gain due to the fusion process. This indicates that zero-shot score matching can be used as an effective and computationally inexpensive tool to accurately perform data fusion from data of varying sources and spatiotemporal fidelity. Based on the promising results from this study, future extensions of this framework will include the integration of additional dynamical considerations for diffusion-based forecasting, i.e., performing lead-time-dependent score-matching with a diffusion model that generates trajectories, integrating both Eulerian and Lagrangian observations within a common fusion paradigm, and optimal sensor placement for high-quality state recovery.
@@ Acknowledgments
This research used resources of the Argonne Leadership Computing Facility, which is a U.S. Department of Energy Office of Science User Facility operated under contract DE-AC02-06CH11357. DC, HG, and RM acknowledge support from DOE ASCR award "Inertial neural surrogates for stable dynamical prediction". DC, HG, RM also acknowledge the support of computing resources through an AI For Science allocation at the National Energy Research Scientific Computing Center (NERSC) and from Penn State Institute for Computational and Data Sciences (ICDS). RM acknowledges the support of a YIP from ARO Modeling of Complex Systems (PM: Robert Martin).